# A note on the Maxwell’s eigenvalues on thin sets

### v1

@@ Abstract.
We analyse the Maxwell’s spectrum on thin tubular neighborhoods of embedded surfaces of $\mathbb{R}^{3}$. We show that the Maxwell eigenvalues converge to the Laplacian eigenvalues of the surface as the thin parameter tends to zero. To achieve this, we reformulate the problem in terms of the spectrum of the Hodge Laplacian with relative conditions acting on co-closed differential $1$-forms. The result leads to new examples of domains where the Faber-Krahn inequality for Maxwell’s eigenvalues fails, examples of domains with any number of arbitrarily small eigenvalues, and underlines the failure of spectral stability under singular perturbations changing the topology of the domain. Additionally, we explicitly produce the Maxwell’s eigenfunctions on product domains with the product metric, extending previous constructions valid in the Euclidean case.
@@ 1. Introduction and statement of the main results
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$. The second-order reformulation of the time-harmonic Maxwell system
@@
where $\nu$ is the outer normal to $\partial\Omega$ and $\lambda:=\eta^{2}$ is the eigenvalue. If $\Omega$ is sufficiently regular, e.g., if $\partial\Omega$ is Lipschitz, it is well-known that problem (1.1) has $\lambda=0$ as eigenvalue of infinite multiplicity and it further admits a sequence of non-negative eigenvalues of finite multiplicity
@@
We are mainly interested in the dependence of $\lambda_{j}(\Omega)$ upon perturbation of the domain $\Omega$; in particular, we will assume that $\Omega$ is a thin domain, described by tubes of size $h$ around smooth embedded surfaces (with or without boundary). More precisely, if $\Sigma$ is a smooth embedded orientable compact surface in $\mathbb{R}^{3}$ (with or without boundary), we define, for all $h>0$ sufficiently small, the tube $\Omega_{h}$ by
@@
In the case that the boundary is just Lipschitz, problem (1.1), and in particular, the boundary condition $\nu\times E|_{\partial\Omega_{h}}=0$, has to be interpreted in a suitable weak sense, see [11]. We now state our main result.
@@
An immediate consequence of Theorem 1.1 is that we can always find examples of domains $\Omega$ with any number of arbitrary small eigenvalues $\lambda_{j}(\Omega)$, in the class of domains with prescribed volume $|\Omega|$ (see [3] for a recent related result).
@@
Let us first prove the result in the case of $\Omega$ not homeomorphic to a ball. Let $\Sigma$ be a Cheeger’s dumbbell [13, p. 79] with N1N-1 thin passages such that $\mu_{1},...,\mu_{N}<\frac{\epsilon}{2}$ (in particular, $\mu_{1}=0$). Choose $h>0$ so small that the following inequality holds for $\lambda_{j}(\Omega_{h})<\epsilon$ for j=1,Nj=1,...N. Now, for $h$ small, $|\Omega_{h}|\approx|\Sigma|h$ and we can choose $h$ such that $|\Omega_{h}|<1$. Then take $\Omega:=\frac{\Omega_{h}}{|\Omega_{h}|^{1/3}}$ so that $|\Omega|=1$. Then $\lambda_{j}(\Omega)=|\Omega_{h}|^{2/3}\lambda_{j}(\Omega_{h})<\epsilon$ for all j=1,,Nj=1,...,N.
To produce a domain homeomophic to a ball, replace $\Sigma$ in the construction above with $\Sigma_{\delta}:=\Sigma\setminus B_{\delta}$, where $B_{\delta}\subset\Sigma$ is a geodesic disk with $\delta>0$ small enough so that the Dirichlet eigenvalues $\mu_{j}^{D}$ on $\Sigma_{\delta}$ satisfy $\mu_{1}^{D},...,\mu_{N}^{D}<\frac{2\epsilon}{3}$. In fact, as $\delta\to 0^{+}$, the Dirichlet spectrum of $\Sigma_{\delta}$ converges to the spectrum of the Laplacian on $\Sigma$, see e.g., [16]. The rest of the construction is as in the previous part of the proof. Just note that a thin tube around $\Sigma_{\delta}$ (or, equivalently, $\Sigma_{\delta}\times(0,h)$) is homeomorphic to a ball. ∎
Corollary 1.2 implies that a Faber-Krahn inequality cannot hold for the first Maxwell eigenvalue, nor for other functions of the eigenvalues like the sum or the product (or other elementary symmetric functions) of the first $N$ eigenvalues, as already highlighted in [29] (see also [40]).
@@
where now $u$ is a differential $1$-form on $M$, $\Delta=d\delta+\delta d$ is the Hodge Laplacian associated with the metric $g$ acting on differential forms, $d$ is the exterior derivative, $\delta$ is the codifferential associated with the metric $g$ and $i:\partial M\to M$ is the canonical inclusion. We refer to Section 2 for more details. When $M$ is a bounded domain in $\mathbb{R}^{3}$ and the metric is the Euclidean one, problems (1.1) and (1.3) coincide (under the canonical identification of vector fields and $1$-forms).
A useful observation is to realise that for $h$ small, the domain $\Omega_{h}$ with the Euclidean metric is quasi-isometric to the manifold $M=\Sigma\times(0,h)$ with the product metric $g_{p}=g_{\Sigma}\times dt^{2}$, where $g_{\Sigma}$ is the induced metric on $\Sigma$ from the ambient Euclidean space. Using the product structure of the metric, we are able to explicitly describe all the eigenvalues of problem (1.3) on $M=\Sigma\times(0,h)$ and the associated eigenfunctions. Concerning the eigenvalues, we have the following (see Theorems 3.1 and 3.4)
@@
Let $M=\Sigma\times(0,h)$, $h>0$, $(\Sigma,g_{\Sigma})$ be a compact Riemannian surface (without boundary) and $g_{p}=g_{\Sigma}+dt^{2}$ be the product metric on $M$. Then the spectrum of (1.3) on $(M,g_{p})$ is given by the union the following four families:
@@
Here $\mu_{k}^{D},\mu_{k}^{N}$ are the eigenvalues of the Laplacian on $(\Sigma,g_{\Sigma})$ with Dirichlet and Neumann boundary conditions, respectively (with multiplicities), $\eta_{j}(h),d_{j}(h)$ are the Neumann and Dirichlet eigenvalues on $(0,h)$, $\gamma$ is the genus of the surface and $b+1$ is the number of connected components of $\partial\Sigma$.
@@
Theorems 1.3 and 1.4 say that the TE-TM-TEM description of the eigenvalues given in [15] continues to hold in the Riemannian setting, that is, when we replace $\omega$ with a Riemannian surface $\Sigma$, generalising therefore the construction valid for straight cylinders to possibly curved ones. The only difference is that the Maxwell eigenvalues will now be described in terms of the eigenvalues of the Laplacian on the surface $\Sigma$ (with Dirichlet or Neumann boundary conditions on $\partial\Sigma$ when $\partial\Sigma\neq\emptyset$, as in the flat case).
From the description in [15], we easily see that the limiting spectrum of the Maxwell operator on the flat cylinder $\omega\times I$ as $|I|\to 0^{+}$ coincides with the Dirichlet spectrum of the Laplacian on $\omega$: this is a particular case of Theorem 1.1.
Another implication of Theorem 1.1 is that there is no spectral stability under singular domain perturbation when a change of topology is involved. More precisely, let $\Sigma_{\delta}=\omega\setminus B_{\delta}\subset\mathbb{R}^{2}$, where $\omega$ is a simply connected planar domain, $B_{\delta}$ is a disk of radius $\delta$ centered at some $x\in\omega$, $B_{\delta}\subset\omega$ for all $\delta>0$ sufficiently small. Let $\Omega_{h,\delta}=\Sigma_{\delta}\times(0,h)$. Let $h>0$ be fixed. Then the Maxwell eigenvalues on $\Omega_{h,\delta}$ are just those given by Theorem 1.4:
@@
When $\delta\to 0^{+}$, $\Omega_{h,\delta}$ converges (in the sense of Hausdorff convergence) to $\Omega_{h}=\omega\times(0,h)$. The Maxwell spectrum on $\Omega_{h}$ is given by Theorem 1.4; however, we note that, since $\omega$ is simply connected, we do not have the third family of eigenvalues:
@@
The first two families of eigenvalues behave continuously in $\delta$ (this follows from the spectral stability of the Dirichlet and Neumann eigenvalues on Euclidean domains under removal of a small ball). On the contrary, the eigenvalues of the third family clearly admit a limit as $\delta\to 0^{+}$ (they do not depend on $\delta$), but the limits are not Maxwell’s eigenvalues on the limit domain.
@@
Associated with the families of eigenvalues in Theorems 1.3 and 1.4, there are families of eigenfunctions that we describe more explicitly in Theorems 3.1 and 3.4. In these theorems the eigenfunctions are interpreted as eigenfunctions of the Hodge Laplacian (problem 3.1), that is, they are $1$-forms. Finally, we prove that the eigenfunctions on $\Omega_{h}$ and the limit eigenfunctions on $\Sigma$ converge in a suitable sense as $h\to 0^{+}$, see Theorem 5.1 and Corollary 5.2.
@@
The analysis of eigenvalue problems for differential operators on thin domains is a classical topic that has experienced a noticeable growth in recent years, see e.g., [4, 5, 7, 8, 26, 30, 31] and references therein. Our analysis was inspired by the well-known result in [41]: the Neumann eigenvalues of a thin tube around a closed embedded hypersurface in $\mathbb{R}^{n}$ converge to the eigenvalues of the Laplacian on the surface. Since then, the analysis of the behavior of the spectrum in the thin limit turned out to be useful in the study of many other spectral problems, e.g., the hot spot conjecture [32], the clamped plate equation [12], Navier-Stokes equations [35, 34, 36], quantum waveguides [17, 18, 19, 39]. From the geometric point of view, the analysis of the spectrum of the Hodge Laplacian acting on $p$-forms on domains with thin parts has been considered by various authors, see e.g., [1, 2, 3]. However, also from the geometric point of view, we were not aware of a result in the spirit of [41] for $p$-forms. This is the main motivation of the present note, which focuses on $p=1$ due to the relation of the problem on forms with the Maxwell’s problem. We finally remark that in the last ten years there has been an upsurge of interest in the connection between the spectrum of the Maxwell operator and the underlying geometry, mainly in the Euclidean setting, see for instance [6, 9, 20, 21, 22], where, for instance, the role of the topology of the domain in the spectral properties of the Maxwell system prominently appears.
@@
The present note is organised as follows. Section 2 contains a few geometric preliminaries and the description of the connection between problems (1.1) and (1.3). In Section 3 we prove Theorems 1.3 and 1.4, namely, we describe the Maxwell eigenvalues and the eigenfunction of the product manifold $(M,g_{p})=(\Sigma\times(0,L),g_{p})$. In Section 4 we prove our main Theorem 1.1. In Section 5 we establish a convergence result for eigenfunctions. Finally, we remark that in this note we identify the Maxwell problem with the eigenvalue problem for the Hodge Laplacian restricted on co-closed $1$-forms with relative conditions. For the reader’s convenience, in Section A we describe the spectrum of the full Hodge Laplacian with relative boundary conditions on the product manifold $(M,g_{p})$.
@@
We conclude this section by underlying that this article is purposely written to be accessible to both mathematical analysts and differential/spectral geometers, and therefore it may contain details that are usually omitted in a research article.
@@ 2.1. Notation and functional spaces
is compactly embedded in $L^{2}(\Omega)^{3}$. Let us now consider the weak fomulation of equation (1.1), that is
@@
There are now two (equivalent) ways of studying the spectrum of this operator. Either we study it in the Hilbert space $L^{2}(\Omega)^{3}$, and then $\lambda=0$ is a point of essential spectrum of the operator; or we restrict the Hilbert space to $H(\operatorname{div}0,\Omega)=\{E\in L^{2}(\Omega)^{3}:\operatorname{div}E=0\}$, which corresponds to restrict the domain of the operator $\operatorname{curl}\operatorname{curl}$ to the orthogonal of its (infinite-dimensional) kernel. We proceed with this second option.
@@
for all $u\in\operatorname{dom}(T):=\{u\in\operatorname{dom}(Q):T^{1/2}u\in\operatorname{dom}(Q)\}$, $v\in\operatorname{dom}(Q)$. Moreover, the compact embedding of $\operatorname{dom}(Q)$ into the ambient Hilbert space $H(\operatorname{div}0,\Omega)$ implies that the resolvent $T^{-1}$ is compact as an operator in $H(\operatorname{div}0,\Omega)$. By standard spectral theory we deduce that the spectrum of $T$ coincides with its discrete spectrum, and can be described by a sequence of positive, isolated eigenvalues of finite multiplicity
$0<\lambda_{1}(\Omega)\leq\lambda_{2}(\Omega)\leq\dots\leq\lambda_{j}(\Omega)\leq\dots\nearrow+\infty,$

### v2

@@ Abstract.
We analyse the Maxwell’s spectrum on thin tubular neighborhoods of embedded surfaces of $\mathbb{R}^{3}$. We show that the Maxwell’s eigenvalues converge to the Laplacian eigenvalues of the surface as the thin parameter tends to zero. To achieve this, we reformulate the problem in terms of the spectrum of the Hodge Laplacian with relative conditions acting on co-closed differential $1$-forms. The result leads to new examples of domains where the Faber-Krahn inequality for Maxwell’s eigenvalues fails, examples of domains with any number of arbitrarily small eigenvalues, and underlines the failure of spectral stability under singular perturbations changing the topology of the domain. Additionally, we explicitly produce Maxwell’s eigenfunctions on product domains with the product metric, extending previous constructions valid in the Euclidean case.
@@ 1. Introduction and statement of the main results
Let $\Omega$ be a bounded domain in $\mathbb{R}^{3}$. The second-order reformulation of the time-harmonic Maxwell’s system
@@
where $\nu$ is the outer normal to $\partial\Omega$ and $\lambda:=\eta^{2}$ is the eigenvalue. If $\Omega$ is sufficiently regular, e.g., if $\partial\Omega$ is Lipschitz, it is well-known that problem (1.1) has $\lambda=0$ as eigenvalue of infinite multiplicity and it further admits a sequence of non-negative eigenvalues of finite multiplicity (corresponding to divergence-free eigenfunctions)
@@
We are mainly interested in the dependence of $\lambda_{j}(\Omega)$ upon the perturbation of the domain $\Omega$; in particular, we will assume that $\Omega$ is a thin domain, described by tubes of size $h$ around smooth embedded surfaces (with or without boundary). More precisely, if $\Sigma$ is a smooth, embedded, orientable, compact surface in $\mathbb{R}^{3}$ (with or without boundary), we define, for all $h>0$ sufficiently small, the tube $\Omega_{h}$ by
@@
If the boundary is just Lipschitz, problem (1.1), and in particular, the boundary condition $\nu\times E|_{\partial\Omega_{h}}=0$, has to be interpreted in a suitable weak sense, see [11]. We now state our main result.
@@
An immediate consequence of Theorem 1.1 is that we can always find examples of domains $\Omega$ with any number of arbitrarily small eigenvalues $\lambda_{j}(\Omega)$, in the class of domains with prescribed volume $|\Omega|$ (see [3] for a recent related result).
@@
Let us first prove the result in the case of $\Omega$ not homeomorphic to a ball. Let $\Sigma$ be a Cheeger’s dumbbell [13, p. 79] with N1N-1 thin passages such that $\mu_{1},...,\mu_{N}<\frac{\epsilon}{2}$ (in particular, $\mu_{1}=0$). Choose a small enough $h>0$ guaranteeing that $\lambda_{j}(\Omega_{h})<\epsilon$ for j=1,Nj=1,...N. Now, for $h$ small, $|\Omega_{h}|\approx|\Sigma|h$ and we can possibly reduce $h$ in such a way that $|\Omega_{h}|<1$. Let $\Omega:=\frac{\Omega_{h}}{|\Omega_{h}|^{1/3}}$, so that $|\Omega|=1$. Then $\lambda_{j}(\Omega)=|\Omega_{h}|^{2/3}\lambda_{j}(\Omega_{h})<\epsilon$ for all j=1,,Nj=1,...,N.
To produce a domain homeomorphic to a ball, replace $\Sigma$ in the above construction with $\Sigma_{\delta}:=\Sigma\setminus B_{\delta}$, where $B_{\delta}\subset\Sigma$ is a small geodesic disk with $\delta>0$, chosen in such a way that the Dirichlet eigenvalues $\mu_{j}^{D}$ in $\Sigma_{\delta}$ satisfy $\mu_{1}^{D},...,\mu_{N}^{D}<\frac{2\epsilon}{3}$. This is possible because the Dirichlet spectrum of $\Sigma_{\delta}$ converges to the spectrum of the Laplacian on $\Sigma$ as $\delta\to 0^{+}$, see e.g., [16]. The rest of the construction is as in the previous part of the proof. It is sufficient to note that a thin tube around $\Sigma_{\delta}$ (or, equivalently, $\Sigma_{\delta}\times(0,h)$) is homeomorphic to a ball. ∎
Corollary 1.2 implies that a Faber-Krahn inequality cannot hold for the first Maxwell’s eigenvalue, nor for other functions of the eigenvalues like the sum or the product (or other elementary symmetric functions) of the first $N$ eigenvalues, as already highlighted in [32] (see also [40]).
@@
where now $u$ is a differential $1$-form on $M$, $\Delta=d\delta+\delta d$ is the Hodge Laplacian associated with the metric $g$ acting on differential forms, $d$ is the exterior derivative, $\delta$ is the codifferential associated with the metric $g$ and $i:\partial M\to M$ is the canonical inclusion. We refer to Section 2 for more details. When $M$ is a bounded domain in $\mathbb{R}^{3}$ and the metric is the Euclidean one, problems (1.1) restricted to divergence-free fields and (1.3) coincide (under the canonical identification of vector fields and $1$-forms).
It is worth recalling the following fact. For small $h>0$, the domain $\Omega_{h}$ with the Euclidean metric is quasi-isometric to the manifold $M=\Sigma\times(0,h)$ with the product metric $g_{p}=g_{\Sigma}\times dt^{2}$, where $g_{\Sigma}$ is the induced metric on $\Sigma$ from the ambient Euclidean space. Using the product structure of the metric, we are able to explicitly describe all the eigenvalues of Problem (1.3) in $M=\Sigma\times(0,h)$ and the associated eigenfunctions. Concerning the eigenvalues, we have the following (see Theorems 3.1 and 3.4)
@@
Let $M=\Sigma\times(0,h)$, $h>0$, $(\Sigma,g_{\Sigma})$ be a compact Riemannian surface (without boundary) and $g_{p}=g_{\Sigma}+dt^{2}$ be the product metric on $M$. Then the spectrum of (1.3) on $(M,g_{p})$ is given by the union of the following four families:
@@
Here $\mu_{k}^{D},\mu_{k}^{N}$ are the eigenvalues of the Laplacian on $(\Sigma,g_{\Sigma})$ with Dirichlet and Neumann boundary conditions, respectively (with multiplicities), $\eta_{j}(h),d_{j}(h)$ are the Neumann and Dirichlet eigenvalues on $(0,h)$, $\gamma$ is the genus of the surface, and $b+1$ is the number of connected components of $\partial\Sigma$.
@@
Theorems 1.3 and 1.4 say that the TE-TM-TEM description of the eigenvalues given in [15] continues to hold in the Riemannian setting, that is, when we replace $\omega$ with a Riemannian surface $\Sigma$, therefore generalising the construction valid for straight cylinders to possibly curved ones. The only difference is that the Maxwell’s eigenvalues will now be described in terms of the eigenvalues of the Laplacian on the surface $\Sigma$ (with Dirichlet or Neumann boundary conditions on $\partial\Sigma$ when $\partial\Sigma\neq\emptyset$, as in the flat case).
From the description in [15], we easily see that the limiting spectrum of the Maxwell’s operator on the flat cylinder $\omega\times I$ as $|I|\to 0^{+}$ coincides with the Dirichlet spectrum of the Laplacian on $\omega$: this is a particular case of Theorem 1.1.
Another implication of Theorem 1.1 is that there is no spectral stability under singular domain perturbation when a change of topology is involved. More precisely, let $\Sigma_{\delta}=\omega\setminus B_{\delta}\subset\mathbb{R}^{2}$, where $\omega$ is a simply connected planar domain, $B_{\delta}$ is a disk of radius $\delta$ centered at some $x\in\omega$, $B_{\delta}\subset\omega$ for all $\delta>0$ sufficiently small. Let $\Omega_{h,\delta}=\Sigma_{\delta}\times(0,h)$. Let $h>0$ be fixed. Then the Maxwell’s eigenvalues on $\Omega_{h,\delta}$ are just those given by Theorem 1.4:
@@
When $\delta\to 0^{+}$, $\Omega_{h,\delta}$ converges (in the sense of Hausdorff convergence) to $\Omega_{h}=\omega\times(0,h)$. The Maxwell’s spectrum on $\Omega_{h}$ is given by Theorem 1.4; however, we note that, since $\omega$ is simply connected, we do not have the third family of eigenvalues:
@@
The first two families of eigenvalues behave continuously in $\delta$ (this follows from the spectral stability of the Dirichlet and Neumann eigenvalues in Euclidean domains upon removal of a small ball). On the contrary, the eigenvalues of the third family clearly admit a limit as $\delta\to 0^{+}$ (they do not depend on $\delta$), but the limits are not Maxwell’s eigenvalues in the limit domain.
@@
Associated with the families of eigenvalues in Theorems 1.3 and 1.4, there are families of eigenfunctions that we describe more explicitly in Theorems 3.1 and 3.4. In these theorems the eigenfunctions are interpreted as eigenfunctions of the Hodge Laplacian (problem 3.1), that is, they are $1$-forms. Finally, we prove that the eigenfunctions in $\Omega_{h}$ and the limit eigenfunctions in $\Sigma$ converge in a suitable sense as $h\to 0^{+}$, see Theorem 5.1 and Corollary 5.2.
@@
The analysis of eigenvalue problems for differential operators on thin domains is a classical topic that has experienced a noticeable growth in recent years, see e.g., [4, 5, 7, 8, 26, 31, 29] and references therein. Our analysis was inspired by the well-known result in [41]: the Neumann eigenvalues of a thin tube around a closed embedded hypersurface in $\mathbb{R}^{n}$ converge to the eigenvalues of the Laplacian on the surface. Since then, the analysis of the behavior of the spectrum in the thin limit turned out to be useful in the study of many other spectral problems, e.g., the hot spot conjecture [30], the clamped plate equation [12], Navier-Stokes equations [35, 34, 36], quantum waveguides [17, 18, 19, 39]. It is worth mentioning the recent paper [LotOB], in which the authors perform an asymptotic analysis of the spectrum of the Dirac operator in tubes around hypersurfaces in $\mathbb{R}^{n}$. When $n=3$, these are the domains considered in the present paper. When the tube shrinks to the surface, they are able to identify the effective Schrödinger operator driving the dynamics, and write an asymptotic expansion of the eigenvalues. From the geometric point of view, the analysis of the spectrum of the Hodge Laplacian acting on $p$-forms on domains with thin parts has been considered by various authors, see e.g., [1, 2, 3]. However, also from the geometric point of view, we were not aware of a result in the spirit of [41] for $p$-forms. This is the main motivation of the present note, which focuses on $p=1$ due to the relation of the problem on forms with the Maxwell’s problem. We finally remark that in the last ten years there has been an upsurge of interest in the connection between the spectrum of the Maxwell’s operator and the underlying geometry, mainly in the Euclidean setting, see, for instance [6, 9, 20, 21, 22], where, for instance, the role of the topology of the domain in the spectral properties of the Maxwell’s system prominently appears.
@@
The present note is organised as follows. Section 2 contains a few geometric preliminaries and the description of the connection between problems (1.1) and (1.3). In Section 3 we prove Theorems 1.3 and 1.4, namely, we describe the Maxwell’s eigenvalues and the eigenfunction of the product manifold $(M,g_{p})=(\Sigma\times(0,L),g_{p})$. In Section 4 we prove our main Theorem 1.1. In Section 5 we establish a convergence result for eigenfunctions. Finally, we remark that in this note we identify the Maxwell’s problem with the eigenvalue problem for the Hodge Laplacian restricted on co-closed $1$-forms with relative conditions. For the reader’s convenience, in Section A we describe the spectrum of the full Hodge Laplacian with relative boundary conditions in the product manifold $(M,g_{p})$.
@@
We conclude this section by underlining that this article is purposely written to be accessible to both mathematical analysts and differential/spectral geometers, and therefore it may contain details that are usually omitted in a research article.
@@ 2.1. Notation and functional spaces
is compactly embedded in $L^{2}(\Omega)^{3}$. Let us now consider the weak formulation of equation (1.1), which reads
@@
There are now two (equivalent) ways of studying the spectrum of this operator. Either we study it in the Hilbert space $L^{2}(\Omega)^{3}$, and then $\lambda=0$ is a point of essential spectrum of the operator; or we restrict the Hilbert space to $H(\operatorname{div}0,\Omega)=\{E\in L^{2}(\Omega)^{3}:\operatorname{div}E=0\}$, which corresponds to restrict the domain of the operator $\operatorname{curl}\operatorname{curl}$ to the orthogonal of $\nabla H_{0}^{1}(\Omega)$. We proceed with this second option.
@@
for all $u\in\operatorname{dom}(T):=\{u\in\operatorname{dom}(Q):T^{1/2}u\in\operatorname{dom}(Q)\}$, $v\in\operatorname{dom}(Q)$. Moreover, the compact embedding of $\operatorname{dom}(Q)$ into the ambient Hilbert space $H(\operatorname{div}0,\Omega)$ implies that the resolvent $T^{-1}$ is compact as an operator in $H(\operatorname{div}0,\Omega)$. By standard spectral theory we deduce that the spectrum of $T$ coincides with its discrete spectrum and can be described by a sequence of non-negative, isolated eigenvalues of finite multiplicity